# Relative Entropy Methods for Calculating Committors

### v1

@@ Appendix A Proofs of Results in Section 3
00 0 $=\nabla_{\theta}\mathds{P}_{\theta}[\Omega]$
$=\nabla_{\theta}\mathds{Q}\left[\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right]$
$=\mathds{Q}\left[\nabla_{\theta}\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right]$
$=\mathds{P}_{\theta}\left[\nabla_{\theta}\log\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right].$ (27)
@@
$\nabla_{\theta}D_{\mathrm{KL}}(\mathds{P}_{\theta}|\mathds{Q})$ $=\mathds{Q}\left[g^{\prime}\left(\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right)\nabla_{\theta}\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right]$
$=\mathds{Q}\left[\left(\log\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}+1\right)\nabla_{\theta}\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right]$
$=\mathds{P}_{\theta}\left[\left(\log\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}+1\right)\nabla_{\theta}\log\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right]$
@@
$=\operatorname{cov}_{\mathds{P}_{\theta}}\left(\log\frac{\nu}{\mu_{\theta}}+\log q_{\theta}(X_{\tau})-\int_{0}^{\tau}\frac{Lq_{\theta}}{q_{\theta}}(X_{s})\,\mathrm{d}s,\right.$
$\qquad\qquad\quad\left.\frac{\nabla_{\theta}q_{\theta}(X_{\tau})}{q_{\theta}(X_{\tau})}-\frac{\nabla_{\theta}\mu_{\theta}}{\mu_{\theta}}-\int_{0}^{\tau}\nabla_{\theta}\frac{Lq_{\theta}}{q_{\theta}}(X_{s})\,\mathrm{d}s\right)$
@@
The fourth equality above follows from (27), and the last follows since $\log\frac{\nu}{\mu_{\theta}}$ and $\frac{\nabla_{\theta}\mu_{\theta}}{\mu_{\theta}}$ are constants that do not depend on $\omega$. ∎
Appendix C Proofs of Results in Section 5
Proof of Lemma 6.
For any $x\in\Omega$, we will write $\bar{x}$ for $\rho(x)$ to simplify notation. Let $w:\Omega\rightarrow\mathbb{R}$ admit a twice continuously differentiable extension to an open neighborhood of $\overline{\Omega}$. Suppose that for some twice continuously differentiable $w_{0}:\partial A\rightarrow\mathbb{R}$, $w_{1}:\partial A\rightarrow\mathbb{R}$, and $w_{2}:\Omega\rightarrow\mathbb{R}$, we have
$w(x)=w_{0}(\bar{x})+T(x)w_{1}(\bar{x})+T(x)^{2}w_{2}(x)$ (28)
for all $x\in\Omega$. We will show that there can exist at most one triple of functions $w_{0}$, $w_{1}$, and $w_{2}$ solving (28). First, observe that if (28) holds, then
$w_{0}(z)=w(z)$ (29)
for all $z\in\partial A$, since $T=0$ on $\partial A$. Thus, given $w$, equation (28) uniquely determines $w_{0}$. To see that (28) also determines $w_{1}$, observe that
$\nabla w(x)^{t}$ $=\nabla w_{0}(x)^{t}D\rho(x)+w_{1}(\bar{x})\nabla T(x)^{t}$
$\qquad+T(x)\Big{\{}\nabla w_{1}(x)^{t}D\rho(x)+2w_{2}(x)\nabla T(x)^{t}\Big{\}}$
$\qquad+T(x)^{2}\nabla w_{2}(x)^{t}.$
Thus, for $z\in\partial A$,
$\nabla w(z)\cdot n(z)=\nabla w_{0}(z)^{t}D\rho(z)n(z)+w_{1}(z)\nabla T(z)\cdot n(z),$ (30)
hence
$w_{1}(z)=\frac{\nabla w(z)\cdot n(z)-\nabla w_{0}(z)^{t}D\rho(z)n(z)}{\nabla T(z)\cdot n(z)},$ (31)
since we assume $\nabla T(z)\cdot n(z)>0$. Finally, we have $T(x)>0$ for $x\in\Omega\setminus\partial A$, so
$w_{2}(x)=\frac{w(x)-w_{0}(\bar{x})-T(x)w_{1}(\bar{x})}{T(x)^{2}}.$ (32)
Thus, equation (28) determines $w_{2}$ over $\Omega\setminus\partial A$, and therefore also on $\partial A$, since we assume $w_{2}$ is continuous.
Now observe that for any given $w$, equations (29), (31), and (32) define smooth $w_{0}$, $w_{1}$, and $w_{2}$ so that (21) holds. To see this, one need only show that $w_{2}$ defined by (32) is smooth even though $T$ is zero on $\partial A$. For
$v(x)=T(x)^{2}w_{2}(x)=w(x)-w_{0}(\bar{x})-T(x)w_{1}(\bar{x}),$
we have
$\nabla v(z)\cdot n(z)=0$
for $z\in\partial A$. Therefore, by an argument similar to the proof of Lemma 2 but carrying the Taylor expansions to higher order,
$w_{2}(x)$ $=\frac{v(x)}{T(x)^{2}}$
is smooth. We conclude that every smooth $w$ admits a unique expression in the form (28).
Finally, we show that $L\tilde{q}=0$ on $\partial A$ if and only if
$w_{1}=-\frac{\nabla w_{0}^{t}D\rho n}{\lvert\nabla T\rvert}-\frac{LT}{\lvert\nabla T\rvert^{2}}.$ (33)
To see this, observe that for $\tilde{q}=T\exp(w)$,
$L\tilde{q}=\exp(w)\Big{\{}LT+2\varepsilon\nabla w\cdot\nabla T+T(Lw+\varepsilon\lvert\nabla w\rvert^{2})\Big{\}}.$
Therefore, since $T=0$ on $\partial A$, $L\tilde{q}=0$ on $\partial A$ if and only if
$LT+2\varepsilon\nabla w\cdot\nabla T=0$ (34)
on $\partial A$. Since $T=0$ and $\nabla T\cdot n>0$ on $\partial A$, $\nabla T=\lvert\nabla T\rvert n$ on $\partial A$. Thus, (34) reduces to the Neumann boundary condition
$\nabla w\cdot n=-\frac{1}{2\varepsilon}\frac{LT}{\lvert\nabla T\rvert}$ (35)
on $\partial A$. Equation (33) follows from this Neumann boundary condition and (30). ∎

### v2

@@ Appendix B Proofs of Results in Section 4
$0$ $=\nabla_{\theta}\mathds{E}_{\mathds{P}_{\theta}}[\Omega]$
$=\nabla_{\theta}\mathds{E}_{\mathds{Q}}\left[\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right]$
$=\mathds{E}_{\mathds{Q}}\left[\nabla_{\theta}\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right]$
(37) $=\mathds{E}_{\mathds{P}_{\theta}}\left[\nabla_{\theta}\log\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right].$
@@
$\nabla_{\theta}D_{\mathrm{KL}}(\mathds{P}_{\theta}|\mathds{Q})$ $=\mathds{E}_{\mathds{Q}}\left[g^{\prime}\left(\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right)\nabla_{\theta}\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right]$
$=\mathds{E}_{\mathds{Q}}\left[\left(\log\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}+1\right)\nabla_{\theta}\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right]$
$=\mathds{E}_{\mathds{P}_{\theta}}\left[\left(\log\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}+1\right)\nabla_{\theta}\log\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right]$
@@
$=\operatorname{cov}_{\mathds{P}_{\theta}}\left(\log\frac{\zeta}{\eta_{\theta}}+\log q_{\theta}(X_{\tau})-\int_{0}^{\tau}\frac{Lq_{\theta}}{q_{\theta}}(X_{s})\,\mathrm{d}s,\right$.
$\qquad\qquad\quad\left.\frac{\nabla_{\theta}q_{\theta}(X_{\tau})}{q_{\theta}(X_{\tau})}-\frac{\nabla_{\theta}\eta_{\theta}}{\eta_{\theta}}-\int_{0}^{\tau}\nabla_{\theta}\frac{Lq_{\theta}}{q_{\theta}}(X_{s})\,\mathrm{d}s\right)$
@@
The fourth equality above follows from (37), and the last follows since $\log\frac{\zeta}{\eta_{\theta}}$ and $\frac{\nabla_{\theta}\eta_{\theta}}{\eta_{\theta}}$ are constants that do not depend on $\omega$.
Appendix C Detailed Description of Numerical Experiments
Here, we explain the numerical experiments outlined in Section 6 in more detail.
C.1 Details related to Section 6.2
We used FENICS to compute the committor function by a finite element method [3]. We minimized the Ritz form (25) with the prescribed boundary conditions over a P1 Lagrange finite element space based on a Delaunay triangulation of a regular rectangular grid of size $256\times 1024$ covering the domain $a,b]\times[-4,4]$.
C.2 Details related to Section 6.4
When calculating an approximate committor by gradient descent, we took as the function $w_{\theta}(x)$ a $4\times 16$ tensor product of cardinal cubic B-splines covering $a,b]\times[-3,3]$. That is, for $\theta\in\mathbb{R}^{4\times 16}$ and $x\in\mathbb{R}^{2}$, we defined
$w_{\theta}(x)=\sum_{i=0}^{3}\sum_{j=0}^{15}\theta_{ij}B_{3}\left(\frac{x_{1}-a}{h_{1}}-i\right)B_{3}\left(\frac{x_{2}-(-3)}{h_{2}}-j\right),$
where
$h_{1}=\frac{b-a}{4-1},\text{ }h_{2}=\frac{3-(-3)}{16-1}$
and $B_{3}$ is the cardinal cubic B-spline
$B_{3}(x):=\frac{1}{6}\begin{cases}(2-\lvert x\rvert)^{3}&\text{ if }1\leq\lvert x\rvert\leq 2\\ 4-6x^{2}-3\lvert x\rvert^{3}&\text{if }\lvert x\rvert\leq 1\\ 0&\text{otherwise}\end{cases}$
For the Adam optimizer, we used the standard parameters listed in [19], except that we took $\epsilon=10^{-4}$ instead of $10^{-8}$ to mitigate Adam’s tendency to make wild changes to some entries of $\theta$ when trajectories visit rare regions far from the transition tube. We took an exponentially decreasing step size schedule
$\eta_{n}=0.1^{1+n/512}.$
Although it is well known that stochastic optimization methods may not converge to a minimizer for such a rapidly decreasing schedule, we obtained better results with an exponentially decreasing schedule than with polynomially decreasing schedules for which convergence is guaranteed. Our batch size was $64$. That is, we calculated $64$ trajectories to estimate $\nabla_{\theta}D_{\mathrm{KL}}(\mathds{P}_{\theta}|\mathds{Q})$ for each Adam step.
To simplify the sampling of initial points of trajectories, we approximated the reactive flux distribution $\tilde{m}$ corresponding to $\tilde{q}$ by a discrete distribution
$m_{d}:=\frac{1}{\eta_{d}}\sum_{i=1}^{N_{d}}\lvert\nabla\tilde{q}(x_{i})\rvert\exp(-U(x_{i})/\varepsilon)\delta_{x_{i}}$
supported on the grid
$x_{i}=\left(a,-3+\frac{6}{N_{d}-1}\right)\text{ for }0\leq i\leq N_{d}-1$
covering the interval $\{a\}\times[-3,3]\subset\partial A$ with $N_{d}=1024$. Here,
$\eta_{d}=\sum_{i=1}^{N_{d}}\lvert\nabla\tilde{q}(x_{i})\rvert\exp(-U(x_{i})/\varepsilon)$
is a normalizing constant.
Finally, for efficiency, instead of computing $q_{1}$ defined in (24) by quadrature at each point of every trajectory biased by an approximate comittor of the form $q_{1}(x_{1})\exp(w_{\theta}(x)(b-x_{1}))$ as in (31), we interpolated a quadrature estimate of $q_{1}$ by a B-spline plus a linear function on a uniform grid of $16$ points covering $[a,b]$. We verified the interpolating spline to be correct to approximately five significant figures.
C.3 Details related to Section 6.5
We used the delta method to produce the standard errors in Tables 1 and 2. That is, to estimate the asymptotic variance of the self-normalized importance sampling estimator
$\bar{\tau}:=\frac{\frac{1}{N}\sum_{k=1}^{N}\exp\left(\int_{0}^{\tau_{k}}\frac{L\tilde{q}}{\tilde{q}}(Y^{k}_{s})\,\mathrm{d}s\right)\tau_{k}}{\frac{1}{N}\sum_{k=1}^{N}\exp\left(\int_{0}^{\tau_{k}}\frac{L\tilde{q}}{\tilde{q}}(Y^{k}_{s})\,\mathrm{d}s\right)},$
of the expected crossover time $\mathds{E}_{\mathds{Q}}[\tau]$, we first calculate the sample averages
$\tilde{\tau}$ $:=\frac{1}{N}\sum_{k=1}^{N}\exp\left(\int_{0}^{\tau_{k}}\frac{L\tilde{q}}{\tilde{q}}(Y^{k}_{s})\,\mathrm{d}s\right)\tau_{k},$
$\bar{Z}$ $:=\frac{1}{N}\sum_{k=1}^{N}\exp\left(\int_{0}^{\tau_{k}}\frac{L\tilde{q}}{\tilde{q}}(Y^{k}_{s})\,\mathrm{d}s\right),$
and the covariance matrix
$\bar{\Sigma}:=\begin{pmatrix}\Sigma_{\tilde{\tau},\tilde{\tau}}&\Sigma_{\tilde{\tau},\bar{Z}}\\ \Sigma_{\bar{Z},\tilde{\tau}}&\Sigma_{\bar{Z},\bar{Z}}\\ \end{pmatrix}$,
where $\Sigma_{\tilde{\tau},\bar{Z}}$ is the empirical covariance
$\Sigma_{\tilde{\tau},\bar{Z}}:=\frac{1}{N-1}\sum_{k=1}^{N}\left\{\exp\left(\int_{0}^{\tau_{k}}\frac{L\tilde{q}}{\tilde{q}}(Y^{k}_{s})\,\mathrm{d}s\right)\tau_{k}-\tilde{\tau}\right\}\left\{\exp\left(\int_{0}^{\tau_{k}}\frac{L\tilde{q}}{\tilde{q}}(Y^{k}_{s})\,\mathrm{d}s\right)-\bar{Z}\right\},$
of $\tilde{\tau}$ with $\bar{Z}$ and the other entries of $\bar{\Sigma}$ are defined similarly. We then define $g(x,y)=x/y$ so that $\bar{\tau}=g(\tilde{\tau},\bar{Z})$, and the delta method [5] yields
$s^{2}:=\frac{1}{N}\nabla g(\tilde{\tau},\bar{Z})^{\text{t}}\bar{\Sigma}\nabla g(\tilde{\tau},\bar{Z})=\frac{1}{N}\begin{pmatrix}1/\bar{Z},&-\tilde{\tau}/\bar{Z}^{2}\end{pmatrix}\bar{\Sigma}\begin{pmatrix}1/\bar{Z}\\ -\tilde{\tau}/\bar{Z}^{2}\end{pmatrix}$.
as an estimate of the asymptotic variance of $\bar{\tau}$.